A stone dropped from a certain height can reach the ground in 5s. It is stopped after 3 seconds of its fall and then allowed to fall again. Find the time taken by the stone to reach the ground for the remaining distance.

Grade:11

4 Answers

Hamza
7 years ago
S=ut+1/2at×t1/2+a×t×t125mts so height from which object is dropped is 125mts125/5×3=45mtsDistance left to be traveled is 125-45=80mtsTherefore time taken=√(2s/a)=4seconds ansSince uis the initial velocity it is zero in all the cases
vinay sharma
7 years ago
remaining height =1/2g5x5−1/2g3x3  = 125−45=80
time taken =80=1/2gtxt
txt=16
t=4
hence time taken for remaining journey=4 sec
 
Zubair
6 years ago
First find the total height covered by the stone in 5seconds.
Which can be calculated as
h=Vit+0.5gt2
since stone is dropped its initial velocity will be 0
therefore
h=0.5gt2
since t=5
H=122.5m is the total height covered by the stone in 5sec.
Now find the distance covered in 3sec by using procedure:
h=0.5(9.8)(3)2=44.1m is the distance covered in 3sec.
now subtract Distance From Height you will get remaining distance which is to be covered
122.5-44.1=78.4m(now we want the time for this remaining distance)
Since h=0.5gt
T=(2h/g)1/2
T={2(78.4)/9.8}0.5=4sec
Hope You Will Understand.
shifa
5 years ago
Because parachute has a more surface area than the stone
this is the easy answer i guess 
i hope you understood